# GEMINUS: Dual-aware Global and Scene-Adaptive Mixture-of-Experts for End-to-End Autonomous Driving

### v3

@@ III-D Uncertainty-aware Routing Mechanism
The raw input $x$ is first processed by encoders to form the feature vector $F$. The router then computes a probability distribution over experts from the feature vector $F$, denoted as $P(x)=[p_{1},p_{2},...,p_{N}]$. Subsequently, the Information Entropy[33] of this distribution is calculated to reflect the uncertainty of the router’s decision:
$H(P(x))=-\sum_{i=1}^{N}p_{i}\log(p_{i})$ (4)
To normalize this entropy to a $[0,1]$ range, it is divided by the theoretical maximum entropy. This maximum occurs when probabilities are uniformly distributed across all experts ( $i.e.,pi=1/N\,for\,all\,i$), and its value is $log(N)$. Thus, the Normalized Information Entropy $U(x)$ is defined as:
@@
This Normalized Information Entropy $U(x)$ serves as the measure of scenario uncertainty. A value close to 0 indicates high certainty, meaning the scenario is distinct and the router is confident. Conversely, a value close to 1 indicates high uncertainty meaning the scenario is ambiguous and the router is undecided.
@@ III-E Loss Design
Global Expert Loss. The Global Expert aims to provide a robust, generalized driving policy. Its loss $\mathcal{L}_{Global}$ comprises three main terms.
Trajectory Imitation Loss. This term encourages the Global Expert to accurately predict future waypoints. It minimizes the L1 distance between predicted and ground truth waypoints, formalized as:
$\mathcal{L}_{\text{traj\_global}}=\sum_{t=1}^{K}\left\|w_{t}-\hat{w}_{t}\right\|_{1}$ (6)
Where $w_{t}$ and $\hat{w}_{t}$ are the ground truth and predicted waypoints, respectively, at step $t$ within a prediction horizon of $K$ steps.
Feature Alignment Loss. This loss ensures consistent feature representation. It measures the L2 distance between the Global Expert’s output features and the corresponding expert feature. This serves as an additional supervision signal [18], formalized as:
$\mathcal{L}_{F\_global}=\left\|j_{\text{global}}-j_{\text{expert}}\right\|_{2}$ (7)
The $j_{global}$ denotes the intermediate feature representation from the Global Expert, and $j_{expert}$ is the corresponding feature from the expert for alignment.
Value Alignment Loss. This term guides the Global Expert to predict the expected return for the current state. It employs an L2 loss, formalized as:
$\mathcal{L}_{V\_global}=\left\|v_{\text{global}}-v_{\text{expert}}\right\|_{2}^{2}$ (8)
The $v_{global}$ is the value predicted by the Global Expert’s value branch, and $v_{expert}$ is the corresponding value from Think2Drive[18] expert.
Global Expert Loss is combined as:
@@
$\lambda_{traj}$, $\lambda_{F}$, and $\lambda_{V}$ are tunable loss weights.
Scene-Adaptive Experts Group Loss. The Scene-Adaptive Experts Group comprises $N$ distinct experts. Each expert is trained to master policies for specific scenarios. The loss for this group, $L_{Adaptive}$, is computed as a weighted sum of individual expert losses. For a given sample $x$, only the adaptive expert it is routed to contributes to the loss. If $x$ is routed to expert $e_{i}$, its loss is calculated. This calculation is similar to the components of the Global Expert Loss. It is formalized as:
$\begin{split}\mathcal{L}_{\text{Adaptive}}&=\mathbb{1}(x\rightarrow e_{k})\cdot\Big{(}\lambda_{\text{traj}}\,\mathcal{L}_{\text{traj},k}(x)\\ &\quad+\lambda_{F}\,\mathcal{L}_{F,k}(x)+\lambda_{V}\,\mathcal{L}_{V,k}(x)\Big{)}\end{split}$ (10)
For integer $k\in{1,...,N}$, $\mathcal{L}_{traj,k}(x)$, $\mathcal{L}_{F,k}(x)$, and $\mathcal{L}_{V,k}(x)$ are the trajectory imitation, feature alignment, and value prediction losses for expert $e_{k}$ on sample $x$. These are similar to those defined for the Global Expert. $1(\cdot)$ is the indicator function, ensuring that only the activated expert’s loss contributes for that specific sample.
Router Loss. The Router Loss is designed to effectively train the Dual-aware Router to make accurate expert selection decisions. This loss corresponds to $\mathcal{L}_{\text{scenario}}$, as described in Equation (3).
Speed Prediction Loss. To enhance the agent’s ability to estimate its current dynamic state, a dedicated speed prediction head is integrated. This head predicts the current ego-vehicle speed from the encoded feature. An L1 loss is employed for this prediction. It minimizes the absolute difference between the predicted and ground truth speeds, denoted as $L_{speed}$.
Total Loss. The total loss function for training the GEMINUS model is the weighted sum of all aforementioned loss components:
@@
The $\lambda_{Global}$, $\lambda_{Adaptive}$, $\lambda_{scenario}$, and $\lambda_{speed}$ are empirically determined weighting coefficients. They balance the contributions of each loss term.
@@
Avg. L2 (m)$\downarrow$ Driving Score $\uparrow$ Success Rate(%) $\uparrow$
@@
Avg. L2 is averaged over the predictions in 2 seconds under 2Hz. * denotes expert feature distillation.
@@
TABLE III: Ablation Study on Bench2Drive Benchmark
@@
GEMINUS 65.39 37.73 37.77
ScenarioMoE-E2E (w/o 1) $62.38$ $32.27$ $34.46$
VanillaMoE-E2E (w/o 1+ 2) $59.23$ $29.09$ $32.05$
SingleExpert-E2E (w/o 1+ 2+ 3) $60.73$ $30.91$ $31.63$
For consistency, VanillaMoE-E2E employs five experts with Top-1 sparse activation. 1 denotes uncertainty-aware routing and Global Expert. 2 denotes scenario-aware routing. 3 denotes Mixture-of-Experts.
TABLE IV: Router Accuracy In Different Scenarios
@@
TABLE V: Expert Utilization In Different Scenarios
@@ IV-A Experimental Setup
Implementation Details. The resolution of the input RGB image is 900 $\times$ 256 pixels. The future prediction steps are set to $K$ = 4, with a prediction frequency of 2 Hz. For PID control settings, we adopt the well-tuned parameters proposed in Transfuser [4]. Specifically, for longitudinal control, the PID parameters are set to $K_{P}$ = 5.0, $K_{I}$ = 0.5, and $K_{D}$ =1.0. For lateral control, the PID parameters are $K_{P}$ = 0.75, $K_{I}$ = 0.75, and $K_{D}$ = 0.3. As for the uncertainty threshold $\tau$, the optimal value $\tau$ = 0.5 is determined through experiments. During the training phase, the training dataset is initially divided into five major scenario subsets. This division is inspired by Bench2Drive’s classification of driving skills. Each subset contains samples with specific scenario ID. For the Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign subsets, the respective scenario IDs are $[0,1,2,3,4]$. The loss function weight coefficients are configured as follows: $\lambda_{traj}$ = 1, $\lambda_{F}$ = 0.05, $\lambda_{V}$ = 0.001, $\lambda_{Global}$ = 1, $\lambda_{Adaptive}$ = 1, $\lambda_{scenario}$ = 1, and $\lambda_{speed}$ = 0.05. For all experiments, the model is trained on a single NVIDIA GeForce RTX 4090 GPU using a batch size of 96 for 32 epochs. The Adam optimizer is employed with an initial learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-7}$. The learning rate is reduced by a factor of 2 after 30 epochs.
@@ IV-B Comparison with State-of-the-Art works
As shown in Table I, GEMINUS achieves state-of-the-art performance on Bench2Drive closed-loop benchmark for both Driving Score and Success Rate. Notably, GEMINUS relies solely on monocular visual input, and surpasses existing methods on the Bench2Drive benchmark that use 6-camera images inputs.
While GEMINUS does not exhibit superior performance in open-loop average L2 error, such metrics primarily indicate model convergence rather than reliably assessing real-world driving. In contrast, closed-loop metrics offer a more robust evaluation of actual driving capabilities, a point emphasized by previous research such as TransFuser++ [34] and Bench2Drive [32].
When focusing solely on monocular vision methods, GEMINUS significantly improves upon existing state-of-the-art monocular vision method–TCP-traj*[11]. GEMINUS achieves a 9.17% increase in Driving Score, a 25.77% increase in Success Rate, and a reduction of 5.88% in open-loop average L2 error. Furthermore, as shown in Table II, a 10.37% increase in MultiAbility-Mean.
@@ IV-C Ablation Study
Comparing VanillaMoE-E2E with SingleExpert-E2E. It is obvious that directly introducing a generic MoE framework which is commonly used in LLMs into autonomous driving does not improve model performance. Without specific adaptation, it even leads to a slight decrease in Driving Score and Success Rate. This substantiates our hypothesis: end-to-end autonomous driving systems demand a more tailored MoE framework. Such a framework should specifically address the diverse and complex nature of real-world driving scenarios.
@@
Comparing GEMINUS with ScenarioMoE-E2E. Further incorporating the uncertainty-aware routing mechanism and the Global Expert yields additional performance gains. The Driving Score improved by 4.83%, Success Rate by 22.06%, and MultiAbility-Mean by 19.41%. The integration of the uncertainty-aware routing mechanism and the Global Expert significantly enhances the model’s robustness and stability. This is particularly true in ambiguous scenarios where the router cannot confidently determine the current situation.
@@ IV-D Analysis of Uncertainty Threshold
To investigate the impact of the uncertainty threshold $\tau$ on model performance, the uncertainty threshold $\tau$ is varied from 0.0 to 1.0 with a step size of 0.1, and conduct a series of closed-loop evaluations on the Bench2Drive Benchmark. As depicted in Fig.3, the model’s Driving Score and Success Rate show a trend of initial increase followed by a decrease as $\tau$ gradually increases, reaching its optimum at $\tau$ = 0.5. This indicates that when the router’s uncertainty is less than 0.5, the selection made by the scenario-aware routing is reliable, and the performance of the adaptive experts contributes to improved model performance. Conversely, when the router’s uncertainty is greater than or equal to 0.5, the scenario-aware routing cannot make a reliable decision, necessitating the intervention of the Global Expert to ensure robust and stable performance.
@@ IV-E Router Accuracy and Expert Utilization
Router Accuracy. As depicted in Table IV, the router’s overall scenario prediction accuracy reached 68.06%. It is worth noting that the Traffic Sign subset overlaps with both the Merging and Emergency Brake subsets. In such cases, a single sample might pertain to multiple scenarios. Therefore, the actual prediction accuracy could be even higher. This indicates that the scenario-aware routing can accurately determine the current scenario in most cases. However, it struggles with an accurate prediction in a minority of scenarios. A closer examination of the five validation set scenarios reveals that, in the Overtaking and Traffic Sign scenarios, the router exhibits the highest prediction accuracy. This is mainly because these scenarios have salient visual cues, such as obstacles or traffic signs. These cues significantly enhance the router’s ability to accurately predict the scenario. In contrast, the Give Way scenario presents the lowest prediction accuracy of 2.89%. This discrepancy stems from two primary factors. First, the Give Way subset constitutes only 3.16% of the training set and 4.00% of the validation set. This represents an inherent data imbalance within the official Bench2Drive dataset. Second, GEMINUS relies on monocular visual input. This constrains its ability to detect rear-approaching vehicles in Give Way scenarios, thereby impeding accurate scenario prediction.
@@ V CONCLUSIONS
This paper presents GEMINUS, a novel Dual-aware Mixture-of-Experts framework tailored for end-to-end autonomous driving. Through the effective coupling of the Global Expert and Scene-Adaptive Experts Group via Dual-aware intelligent routing, GEMINUS simultaneously achieves adaptive performance in feature-distinct scenarios and robust performance in feature-ambiguous scenarios. Closed-loop evaluation is conducted on Bench2Drive, GEMINUS outperforms existing methods and achieves state-of-the-art performance in Driving Score and Success Rate, relying solely on monocular visual input. Furthermore, ablation studies demonstrate significant improvements over the original single-expert baseline: 7.67% in Driving Score, 22.06% in Success Rate, and 19.41% in MultiAbility-Mean. In addition, the impact of the uncertainty threshold on model performance is analyzed to determine its optimal value. Furthermore, an in-depth analysis of router accuracy and expert utilization provides insights into GEMINUS’s internal routing mechanism.
This study is limited by the use of monocular camera inputs. To enable the router to consider scene information more comprehensively, the exploration of Dual-aware routing with multi-camera input remains a promising direction for future research. Furthermore, a promising research direction is to replace GEMINUS’s expert networks with Low-rank Adaptation (LoRA) modules, providing a lightweight Mixture-of-Experts plugin that is highly effective for the efficient fine-tuning of pretrained models.

### v4

@@ III-D Uncertainty-aware Routing Mechanism
The router’s predicted probabilities for the N experts, $p_{i}(x)$, as defined in the preceding section, form a distribution $P(x)=[p_{1}(x),p_{2}(x),...,p_{N}(x)]$. Subsequently, the Information Entropy[32] of this distribution is calculated to reflect the uncertainty of the router’s decision:
$H(P(x))=-\sum_{i=1}^{N}p_{i}(x)\log(p_{i}(x))$ (4)
To normalize this entropy to the $[0,1]$ range, it is divided by the theoretical maximum entropy. This maximum occurs when probabilities are uniformly distributed across all experts (i.e., $p_{i}=1/N$ for all $i$), and its value is $log(N)$. Thus, the Normalized Information Entropy $U(x)$ is defined as:
@@
This Normalized Information Entropy $U(x)$ serves as the measure of scenario uncertainty. A value close to 0 indicates high certainty, meaning the scenario is distinct and the router is confident. Conversely, a value close to 1 indicates high uncertainty, meaning that the scenario is ambiguous and the router is undecided.
@@ III-E Loss Design
Global Expert Loss. The Global Expert aims to provide a robust and generalized driving policy. Its loss, $\mathcal{L}_{\text{Global}}$, is formulated as a weighted sum of three distinct components:
@@
where $\mathcal{L}_{\text{traj\_global}}$ is a trajectory imitation loss that minimizes the L2 distance between predicted and ground-truth waypoints. The other two terms, a feature alignment loss $\mathcal{L}_{F\_global}$ and a value alignment loss $\mathcal{L}_{V\_global}$, serve to distill knowledge from the Think2Drive expert [18] by minimizing the L2 distance to the teacher’s intermediate features and predicted value. The $\lambda$ coefficients are tunable loss weights.
Scene-Adaptive Experts Group Loss. The Scene-Adaptive Experts Group comprises $N$ distinct experts. Each expert is trained to master the policies for specific scenarios. The loss for this group, $L_{Adaptive}$, is calculated as a weighted sum of the losses of individual experts. For a given sample $x$, only the adaptive expert to which it is routed contributes to the loss. This calculation is similar to the components of the Global Expert Loss. It is formalized as:
$\begin{split}\mathcal{L}_{\text{Adaptive}}&=\mathbf{1}(x\rightarrow e_{i})\cdot\Big{(}\lambda_{\text{traj}}\,\mathcal{L}_{\text{traj},i}(x)\\ &\quad+\lambda_{F}\,\mathcal{L}_{F,i}(x)+\lambda_{V}\,\mathcal{L}_{V,i}(x)\Big{)}\end{split}$ (7)
For integers $i\in\{1,\dots,N\}$, $\mathcal{L}_{traj,i}(x)$, $\mathcal{L}_{F,i}(x)$, and $\mathcal{L}_{V,i}(x)$ are the trajectory imitation loss, feature alignment loss, and value prediction loss for expert $e_{i}$ on sample $x$. $\mathbf{1}(\cdot)$ is the indicator function that ensures that only the loss of the activated expert contributes to that specific sample.
Total Loss. The total loss function for training the GEMINUS model is the weighted sum of the expert losses, a router loss and an auxiliary speed prediction loss:
@@
Here, $\mathcal{L}_{\text{scenario}}$ is the router’s cross-entropy loss used to train accurate expert selection (as described in Equation (3) ). $\mathcal{L}_{\text{speed}}$ is an L1 loss from a dedicated head that predicts the speed of the ego vehicle to improve its state estimation. The coefficients $\lambda$ are empirically determined weights that balance the contribution of each loss term.
@@
[…]
@@
* denotes expert feature distillation.
Avg.L2 is averaged over the predictions in 2 seconds under 2Hz.
@@
TABLE III: Ablation Study on Bench2Drive Benchmark.
@@
GEMINUS( 1+ 2+ 3) 65.39 37.73 37.77
ScenarioMoE-E2E ( 1+ 2) $62.38$ $32.27$ $34.46$
VanillaMoE-E2E ( 1) $59.23$ $29.09$ $32.05$
SingleExpert-E2E $60.73$ $30.91$ $31.63$
For consistency, VanillaMoE-E2E uses five experts with Top-1 activation.
1
denotes Mixture-of-Experts. 2 denotes scenario-aware routing. 3 denotes uncertainty-aware routing and Global Expert.
TABLE IV: Router Accuracy in Different Scenarios.
@@
TABLE V: Expert Utilization in Different Scenarios.
@@ IV-A Experimental Setup
Implementation Details. The model is trained for 32 epochs on a single NVIDIA GeForce RTX 4090 GPU with a batch size of 96. We use the Adam optimizer with an initial learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-7}$, and halve the learning rate after 30 epochs. The model takes $900\times 256$ pixel RGB images to predict $K=4$ future waypoints at 2 Hz. Inspired by Bench2Drive, the training dataset is divided into five subsets (Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign) with corresponding scenario IDs $[0,1,2,3,4]$. The uncertainty threshold is set to $\tau=0.5$. We adopt PID parameters from Transfuser[5], setting longitudinal ($K_{P},K_{I},K_{D}$) to (5.0, 0.5, 1.0) and lateral to (0.75, 0.75, 0.3). The loss weights are as follows: $\lambda_{traj}=1$, $\lambda_{F}=0.05$, $\lambda_{V}=0.001$, $\lambda_{Global}=1$, $\lambda_{Adaptive}=1$, $\lambda_{scenario}=1$, and $\lambda_{speed}=0.05$.
@@ IV-B Comparison with State-of-the-Art works
As shown in Table I, GEMINUS achieves state-of-the-art performance on Bench2Drive closed-loop benchmark for both Driving Score and Success Rate. Notably, GEMINUS relies solely on monocular visual input, and surpasses existing methods on the Bench2Drive benchmark that use 6 camera.
While GEMINUS does not exhibit superior performance in open-loop average L2 error, such metrics primarily indicate model convergence rather than reliably assess real-world driving. In contrast, closed-loop metrics offer a more robust evaluation of actual driving capabilities, a point emphasized by previous research such as Transfuser++ [38] and Bench2Drive [31].
When focusing solely on monocular vision methods, GEMINUS significantly improves upon existing state-of-the-art monocular vision method, TCP-traj*[11]. GEMINUS achieves a 9.17% increase in Driving Score, a 25.77% increase in Success Rate, and a reduction of 5.88% in open-loop average L2 error. Furthermore, as shown in Table II, a 10.37% increase in MultiAbility-Mean.
@@ IV-C Ablation Study
Comparing VanillaMoE-E2E with SingleExpert-E2E. It is evident that directly introducing a generic MoE framework which is commonly used in LLMs into autonomous driving does not improve model performance. Without specific adaptation, it even leads to a slight decrease in Driving Score and Success Rate. This substantiates our hypothesis: end-to-end autonomous driving systems demand a more tailored MoE framework. Such a framework should specifically address the diverse and complex nature of real-world driving scenarios.
@@
Comparing GEMINUS with ScenarioMoE-E2E. Further incorporating the uncertainty-aware routing mechanism and the Global Expert yields additional performance gains. The Driving Score improved by 4.83%, Success Rate by 16.92%, and MultiAbility-Mean by 9.61%. The integration of the uncertainty-aware routing mechanism and the Global Expert significantly enhances the model’s robustness and stability. This is particularly true in ambiguous scenarios where the router cannot confidently determine the current situation.
@@ IV-D Analysis of Uncertainty Threshold
To investigate the impact of the uncertainty threshold $\tau$ on model performance, the uncertainty threshold $\tau$ is varied from 0.0 to 1.0 with a step size of 0.1 to conduct a series of closed-loop evaluations on the Bench2Drive Benchmark. As depicted in Fig. 4, the Driving Score shows a trend of initial increase followed by a decrease as $\tau$ gradually increases, reaching its optimum at $\tau$ = 0.5. This indicates that when the router’s uncertainty is less than 0.5, the selection made by the scenario-aware routing is reliable, and the performance of the adaptive experts contributes to improved model performance. Conversely, when the router’s uncertainty is greater than or equal to 0.5, the scenario-aware routing cannot make a reliable decision, necessitating the intervention of the Global Expert to ensure robust and stable performance.
Figure 5: Qualitative examples of GEMINUS on the Bench2Drive closed-loop evaluation set. For comparative purposes, while GEMINUS performs closed-loop control, we also visualize the predicted open-loop trajectories from other experts. These include the Global, Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign Expert.
@@ IV-E Router Accuracy and Expert Utilization
Router Accuracy. As depicted in Table IV, the router’s overall scenario prediction accuracy reached 68.06%. It is worth noting that the Traffic Sign subset overlaps with both the Merging and Emergency Brake subsets. In such cases, a single sample pertains to two scenario types. Therefore, the actual prediction accuracy could be even higher. This indicates that the scenario-aware routing can accurately determine the current scenario in most cases. However, it struggles with an accurate prediction in a minority of scenarios. A closer examination of the five validation set scenarios reveals that, in the Overtaking and Traffic Sign scenarios, the router exhibits the highest prediction accuracy. This is mainly because these scenarios have salient visual cues, such as obstacles or traffic signs. These cues significantly enhance the router’s ability to accurately predict the scenario. In contrast, the Give Way scenario presents the lowest prediction accuracy of 2.87%. This discrepancy stems from two primary factors. First, the Give Way subset constitutes only 3.2% of Bench2Drive-base dataset. This represents an inherent data imbalance within the official Bench2Drive dataset. Second, GEMINUS relies on monocular visual input. This constrains its ability to detect rear-approaching vehicles in Give Way scenarios, thereby impeding accurate scenario prediction.
@@
IV-F Qualitative Results
Fig. 5 presents qualitative results for GEMINUS in two complex closed-loop scenarios, showcasing the effectiveness of our Dual-aware Router.
In the signalized left turn scenario (a), when facing a red light, the router correctly selects the Traffic Sign Expert. This choice is critical for safety, as the visualization shows that several other experts predict dangerous trajectories that would proceed illegally through the intersection. In stark contrast, the chosen expert’s trajectory (green) demonstrates correct compliance by stopping safely before the crosswalk. Subsequently, when the light turns green, it executes a smooth and safe turning arc.
Similarly, in the multi-lane merge (b), GEMINUS activates the Merging Expert. The resulting trajectory (yellow) is distinguished by a well-calibrated curvature and the correct directional heading, making it the ideal path for a smooth and effective merge.
@@ V CONCLUSIONS
This paper presents GEMINUS, a novel Dual-aware Mixture-of-Experts framework tailored for end-to-end autonomous driving. Through the effective coupling of the Global Expert and Scene-Adaptive Experts Group via the Dual-aware router, GEMINUS simultaneously achieves adaptive and robust performance in complex and diverse scenarios. Comprehensive evaluations on the Bench2Drive benchmark substantiate the framework’s effectiveness, where GEMINUS achieves state-of-the-art performance and demonstrates significant improvements over baseline methods, relying solely on monocular visual input.
This study is limited by the use of monocular camera inputs. To enable the router to consider scene information more comprehensively, the exploration of Dual-aware Router with multi-camera input remains a promising direction for future research. Additionally, our adherence to the predefined Bench2Drive scenario categories introduces challenges like data imbalance and scenario overlaps, which impede expert specialization and thus limit the full potential of GEMINUS. Future work could focus on designing more rational and fine-grained scenario classification criteria. Furthermore, a promising research direction is to replace GEMINUS’s expert networks with Low-rank Adaptation (LoRA) modules, providing a lightweight Mixture-of-Experts plugin that is highly effective for the efficient fine-tuning of pretrained models.